### v1

@@ 6.3 Shufflecake “Lite”
The idea would be eventually to provide both modes for Shufflecake, “lite” and “full” (the latter including both crash consistency countermeasures and multi-snapshot capabilities), and let the user choose which one is desired during the init operation. In practice, one of the two would be the default choice and the other one could be selected as optional, but deciding which of the two shall be the default and which one optional will require a careful security and usability study. We leave this decision for future work, after proper confrontation with representatives of the envisioned final user demographics.
@@ 6.4 Corruption Resistance
As discussed in Section 6.4, writing data to decoy volumes while not all the hidden volumes are open entails a high risk of corrupting the hidden volumes. There cannot be perfect mitigation to this problem, save for frequent backups. However, it could be possible to reduce this risk of corruption by using some form of error-correction on the unopened volumes (hence sacrificing some space), and then trying to recover the volume if corruption happens. We tested positively this idea using RAID[33], namely by partitioning a Shufflecake hidden volume into different equal-sized logical partitions, and using them to assemble a RAID device. Other methods might be more suitable, such as alpha entanglement codes[14], but the support should be probably “baked-in” into Shufflecake itself in order to be really robust.
@@ 6.5 Use of Disk Metadata
Likewise, one could embed a string specifying a desired mountpoint for that volume. Notice in fact that, given the PD requirements, one cannot let these volumes be assigned at static mountpoints in a regular Linux way, e.g. using fstab or crypttab. Rather, the desired mountpoint should be hidden within the context of the volume itself. Then, if the implementation supports it, once the volume is opened, it can also be automatically mounted at a given position.
Another idea could be to embed virtual quotas, in order to artificially limit the maximum available size of decoy volumes. As it is now, Shufflecake performs maximum overcommitment on the visible available space of all volumes: each of them will appear as large as the underlying device. This can cause issues if the user (or the OS) mistakenly assumes that that space is actually available, and starts writing too much data on the volumes. In order to mitigate this, metadata could be used to limit the size of the block device seen by the OS. Importantly, this must only hold for decoy volumes, because overcommitment is substantially what allows for PD. Therefore, the correct way to implement this is: each volume’s VMB should specify a virtual quota for the volume below itself in the secrecy hierarchy, but not for itself. When an $\ell$-th volume is opened, the virtual quota for all less secret volumes in the hierarchy will be recursively read this way (the lowest volume, e.g. volume $1$, will not have this assigned metadata, or anyway it will be ignored). Then, the $\ell$-th volume could be assigned a virtual size equal to the maximum available size on the device, minus the sum of the virtual quotas of all other volumes. This way, it will always be impossible to accidentally write too much data on the volume hierarchy, but an adversary will always see the most secret unlocked volume as the last one present.
It might also be possible, athough probably overkill and exceedingly complex to implement, to have different security or redundancy policies assigned per-volume rather than per-device, and use the metadata to disambiguate them. For example, it could in theory be possible to have different features in terms of crash consistency (as discussed in Section 6.1), security (Section 6.2) or corruption resistance (Section 6.4) assigned to different volumes within the same hierarchy.
@@ 6.6 Reclaiming Unused Slices
Clearly, we need some sort of hint from the upper layer in order to trigger this operation. To that effect, we need to intercept the trim requests emitted by the file system. These commands are effectively a third instruction accepted by hard disks, besides read and write; they serve as a way for the file system to indicate to the disk that some sectors no longer contain user data, and so the internal disk controller can avoid copying them over when reshuffling its own internal indirection layer [17]. These commands are also vital for the efficiency of disk virtualisation systems, such as Shufflecake, that overcommit the total underlying space and thus need to exploit every occasion to optimise the resource allocation.
Once we have this mechanism in place, we can design a way to reassign a freshly freed-up physical slice to the pool of available ones at a random position, so that the function NewSlice (Algorithm 3) will return it with uniformly random probability when a new slice is required. More concretely: suppose that Shufflecake intercepts an OS signal telling us that a logical slice for volume $V_{i}$ at LSI $\sigma$ is now free. We define a function ReclaimSlice (Algorithm 4) which operates on the same structures used by NewSlice, and also on the position map of the interested volume. This function clears the occupation bitfield of the reclaimed slice and the entry in the position map, then moves the PSI at a random position of prmslices (after octr) by doing another Fisher-Yates iteration. We use a subfunction ReverseShuffle which, given as input a PSI, returns the index of prmslices where this PSI is found, or error if not present. The way to implement this subfunction can vary, e.g. by keeping in-memory a reverse map of prmslices, or by doing a linear search every time.
Algorithm 4 ReclaimSlice($V_{i},\sigma$)
1:$\varPsi\leftarrow\mathtt{GetPosMap}(V_{i},\;\sigma)$
2:$\mathtt{bfld}[\Psi]:=\mathtt{free}$
3:$\mathtt{PosMap}_{i}[\sigma]:=\bot$
4:$k\leftarrow\mathtt{ReverseShuffle}(\Psi)$
5:if $k>\texttt{octr}$ then: return $\triangleright$ No need to reshuffle in this case.
6:$\mathtt{swap}(\mathtt{prmslices}[k],\mathtt{prmslices}[\mathtt{octr}])$
7:$j\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\left[\mathtt{octr},\mathtt{MaxSlices}\right]$
8:$\mathtt{swap}(\mathtt{prmslices}[j],\mathtt{prmslices}[\mathtt{octr}])$
9:if $\mathtt{bfld}[\mathtt{prmslices}[\texttt{octr}]]=\mathtt{free}$ then: $\mathtt{octr}:=\mathtt{octr}-1$
@@ 6.7 Unbounded Number of Volumes
Shufflecake assumes a number max of possible volumes that can be provided by any device. Even if this limit can be chosen freely by implementations, it would be desirable to have a way for creating unlimited numbers of volumes (subject to space availability) per device. This would not only remove an artificial limitation on the scheme, but would also strengthen its operational security by making any kind of safeword-like technique (as discussed in Section 4.2.5) impossible.
Clearly, for this to be achievable, volume headers cannot be adjacent and packed at the beginning of the disk. One idea for further investigation would be to embed every header (except the first, ‘less secret’ one, which is still going to be at the beginning) at random positions within the device, and having them linked by the previous volume header through an ad-hoc pointer field which is always present, and indistinguishable from random without the correct password. Traversing this list of linked headers, however, presents some challenges. In particular, when the user provides one password on volume instantiation, how do we know whether the password is wrong? And, if not, how do we reach the right header unlocked by that password? It might be possible to devise some complex linking scheme for an arbitrary number of “bogus” headers on the device, but in any case the following limitations would apply.
@@
    During init, in case there is a collision during a nxtptr generation over the location of another pre-existing header, the currently generated nxtptr value is discarded and sampled again, until a suitable one is found by brute-forcing (since the total header size is supposed to be negligible in comparison to the device size, this should be very efficient).
@@
The above idea might very well work, but it remains to specify an efficient way to embed this way also the position maps, which might be larger than one slice. Many options are open to evaluation here, from linked lists starting at the header, to multiple branching pointers.
@@ 6.8 Hidden Shufflecake OS
More concretely, eventually Shufflecake could become itself a full PD-focused Linux distribution, where during installation the user is guided in the process of creating volumes and installing other distributions therein in a guided way. For operational efficiency and security, every OS at layer $k$ should be aware of the filesystem and OS in the volume at layers $j<k$ (which is made possible by the hierarchy among Shufflecake volumes). This would also allow a butler daemon to run from the currently running OS and operate in the background on lower-hierarchy OSes, e.g. by performing system updates, downloading emails, etc., so that all these decoy systems are kept up-to-date even if the user neglects to use them regularly. This would in turn allow to ease the suspicion of an adversary when surrendering a decoy password.

### v2

@@ 6.3 Shufflecake “Lite”
The idea would be eventually to provide both modes for Shufflecake, “lite” and “full” (the latter including both crash consistency countermeasures and multi-snapshot capabilities), and let the user choose which one is desired during the $\mathtt{init}$ operation. In practice, one of the two would be the default choice and the other one could be selected as optional, but deciding which of the two shall be the default and which one optional will require a careful security and usability study. We leave this decision for future work, after proper confrontation with representatives of the envisioned final user demographics.
@@ 6.4 Corruption Resistance
As discussed in Section 4.2, writing data to decoy volumes while not all the hidden volumes are open entails a risk of corrupting the hidden volumes. There cannot be perfect mitigation to this problem, save for frequent backups. However, it could be possible to reduce this risk of corruption by using some form of error-correction on the unopened volumes (hence sacrificing some space), and then trying to recover the volume if corruption happens. We tested positively this idea using RAID[33], namely by partitioning a Shufflecake hidden volume into different equal-sized logical partitions, and using them to assemble a RAID device with redundancy. Other methods might be more suitable, such as alpha entanglement codes[14] or other error-correcting codes.
At present, we do not have plans of including corruption-recovery capabilities within Shufflecake itself. This is first because volume corruption is an event which, albeit undesirable, is a consequence of unsafe operations rather than a necessary scenario to be covered. And, second, because we do not want to impose on the user a default choice of mitigation tools (e.g., RAID vs alpha entanglement codes vs something else). However, what we plan of doing is facilitating the integration of such corruption-recovery tools with Shufflecake, by providing the necessary technical machinery to make them work efficiently, and by providing an API interface for communication between Shufflecake, the tools, and the OS.
6.4.1 Improving resolution of corrupted slice assignments.
We have already seen in Section 4.1 how Shufflecake tries to resolve volume corruption by reassigning new slices to disambiguate collisions in the slice maps. We have also seen how this mechanism has the drawback that corruption of a single block of a lower-order volume at the logical level makes the whole slice content unrecoverable to the higher-order volume, because the newly assigned slice contains garbage data. One possible improvement in this respect is to actually clone the corrupted slice instead, so to make (potentially most of) its original content available again to the higher-order volume, thereby greatly improving the effectiveness of external corruption mitigation tools. This has to be done carefully, though: We cannot simply take a fresh available PSI and copy the corrupted physical slice at that position, for two reasons.
The first one is that this would break plausible deniability, not only in the multi-snapshot setting, but also if an adversary gets a single snapshot of the device after a corruption recovery attempt happened by slice cloning as just presented. The reason is that the adversary would see two physical slices equal one-to-one in content, which is clearly unexplainable without assuming that such recovery happened - and hence hinting at the presence of at least two volumes.
The second reason is that the blocks in the cloned slice can only be decrypted (by some of the conflicting corrupted volumes) if the correct IVs are also recovered at the new position. This would indeed happen in the CTR mode of operation we use, because IVs are written explicitly as part of the slice, and would therefore also be copied to the new location, but it would not happen, e.g., in the XTS mode of operation, where IVs depend on the position within the device. This means that such a proposed solution would not be compatible with the ‘lite’ mode of operation of Shufflecake described in Section 6.3 above.
The right approach is therefore to clone the plaintext content of the slice, or at least as much as possible of it, but re-encrypting it (re-randomising it with new IVs) in the new physical slice. This should happen either explicitly (by sampling and storing new fresh IVs on the slice, in the case of CTR mode) or implicitly (by deriving new IVs from the new physical address, in the case of XTS mode). The encryption key to be used, both for decryption and re-encryption, should be that of the higher-order corrupted volume (the one whose position map is being modified). The corrupted blocks within the slice will still be unrecoverable to this volume (but remain available at the original physical position to the lower-order volume which last wrote them), while the non-corrupted blocks will be successfully recovered and re-encrypted to the new location.
The resulting modified $\mathtt{HandleCorruption}$ procedure is depicted in Algorithm 5. This mechanism allows to recover from volume corruption at the block granularity rather than the slice granularity, thereby reducing the overall amount of error on the reconstructed volumes, and increasing the recovery chances of external mitigation tools.
Algorithm 5 $\mathtt{HandleCorruption}(V_{i},\psi,\sigma)$
1:$\psi^{\mathsf{new}}=\mathtt{NewSlice}(V_{i},\sigma)$
2:for $j=0,\ldots,S_{\mathsf{L}}-1$
3:     $\beta_{\mathsf{P}}^{\mathsf{in}}:=\psi\cdot S_{\mathsf{P}}+\Delta_{\mathsf{S}}+j$
4:     $\mathsf{IV}^{\mathsf{in}}\leftarrow\mathtt{LoadIV}\left(\beta_{\mathsf{P}}^{\mathsf{in}}\right)$
5:     ${c}\leftarrow\mathtt{bRead}\left(\beta_{\mathsf{P}}^{\mathsf{in}}\right)$
6:     ${p}:=\mathtt{Decrypt}\left({c},\mathsf{IV}^{\mathsf{in}},\mathsf{VEK}_{i}\right)$
7:     $\beta_{\mathsf{P}}^{\mathsf{out}}:=\psi^{\mathsf{new}}\cdot S_{\mathsf{P}}+\Delta_{\mathsf{S}}+j$
8:     $\mathsf{IV}^{\mathsf{out}}\leftarrow\mathtt{SampleAndStoreIV}\left(\beta_{\mathsf{P}}^{\mathsf{out}}\right)$
9:     ${c}:=\mathtt{Encrypt}\left({p},\mathsf{IV}^{\mathsf{out}},\mathsf{VEK}_{i}\right)$
10:     $\mathtt{bWrite}\left(\beta_{\mathsf{P}}^{\mathsf{out}},{c}\right)$
11:next $j$
12:return
6.4.2 Corruption mitigation API
. Another possible improvement in integrating corruption-mitigation capabilities in Shufflecake is to provide an interface to communicate to upper layers (e.g., the OS, or third-party error-correction tools) whether a volume corruption happened, and possibly where. This can be done, e.g., by publishing a virtual corruption file on /sys/devices/sflc/\$VOLNAME. By detecting whether this file exists, the OS or other recovery tools can know in advance when to start a recovery process without need of user intervention. Whether a volume is corrupted or not should be detected by $\mathtt{dm\mbox{-}sflc}$ during initialisation, as explained above. Then, after the conflicting slice assignment has been resolved, Shufflecake would set a persistent corrupted flag on the affected volume. Whenever a volume is $\mathtt{open}$ed and such a set flag is read, Shufflecake would create the corruption file as explained above. In case extra information on the corruption event is known, for example the LSI of affected slices (only known the first time their corruption is detected, i.e., during the disambiguation procedure explained above), this information could be published, e.g., in /sys/devices/sflc/\$VOLNAME/corruption. This might be helpful in case certain external recovery tools work better by knowing the location of errors.
After the recovery process has completed successfully, the external tool or OS should communicate it to $\mathtt{dm\mbox{-}sflc}$, for example by writing a new file repaired at the same path. Then, on device $\mathtt{close}$, Shufflecake would check if such file exists and, if so, clear the persistent flag on the affected volume.
@@ 6.5 Use of Disk Metadata
Likewise, one could embed a string specifying a desired mountpoint for that volume. Notice in fact that, given the PD requirements, one cannot let these volumes be assigned at static mountpoints in a regular Linux way, e.g. using fstab or crypttab. Rather, the desired mountpoint should be hidden within the context of the volume itself. Then, if the implementation supports it, once the volume is $\mathtt{open}$ed, it can also be automatically mounted at a given position.
Another idea could be to embed virtual quotas, in order to artificially limit the maximum available size of decoy volumes. As it is now, Shufflecake performs maximum overcommitment on the visible available space of all volumes: each of them will appear as large as the underlying device. This can cause issues if the user (or the OS) mistakenly assumes that that space is actually available, and starts writing too much data on the volumes. In order to mitigate this, metadata could be used to limit the size of the block device seen by the OS. Importantly, this must only hold for decoy volumes, because overcommitment is substantially what allows for PD. Therefore, the correct way to implement this is: each volume’s VMB should specify a virtual quota for the volume below itself in the secrecy hierarchy, but not for itself. When an $N$-th volume is $\mathtt{open}$ed, the virtual quota for all less secret volumes in the hierarchy will be recursively read this way (the lowest volume, e.g. volume 00, will not have this assigned metadata, or anyway it will be ignored). Then, the $N$-th volume could be assigned a virtual size equal to the maximum available size on the device, minus the sum of the virtual quotas of all other volumes. This way, it will always be impossible to accidentally write too much data on the volume hierarchy, but an adversary will always see the most secret unlocked volume as the last one present.
Volume metadata is also the natural right place to store a persistent corruption flag, as explained in Section 6.4.
It might also be possible, although probably overkill and exceedingly complex to implement, to have different security or redundancy policies assigned per-volume rather than per-device, and use the metadata to disambiguate them. For example, it could in theory be possible to have different features in terms of crash consistency (as discussed in Section 6.1), security (Section 6.2) or corruption resistance (Section 6.4) assigned to different volumes within the same hierarchy.
@@ 6.6 Reclaiming Unused Slices
Clearly, we need some sort of hint from the upper layer in order to trigger this operation. To that effect, we need to intercept the trim requests emitted by the file system. These commands are effectively a third instruction accepted by hard disks, besides $\mathtt{read}$ and $\mathtt{write}$; they serve as a way for the file system to indicate to the disk that some sectors no longer contain user data, and so the internal disk controller can avoid copying them over when reshuffling its own internal indirection layer [17]. These commands are also vital for the efficiency of disk virtualisation systems, such as Shufflecake, that overcommit the total underlying space and thus need to exploit every occasion to optimise the resource allocation.
Once we have this mechanism in place, we can design a way to reassign a freshly freed-up physical slice to the pool of available ones at a random position, so that the function $\mathtt{NewSlice}$ (Algorithm 3) will return it with uniformly random probability when a new slice is required. More concretely: suppose that Shufflecake intercepts an OS signal telling us that a logical slice for volume $V_{i}$ at LSI $\sigma$ is now free. We define a function $\mathtt{ReclaimSlice}$ (Algorithm 6) which operates on the same structures used by $\mathtt{NewSlice}$, and also on the slice map of the interested volume. This function clears the occupation bitfield of the reclaimed slice and the entry in the slice map, then moves the PSI at a random position of $\mathsf{prmslices}$ (after $\mathsf{octr}$) by doing another Fisher-Yates iteration. We use a subfunction $\mathtt{ReverseShuffle}$ which, given as input a PSI, returns the index of $\mathsf{prmslices}$ where this PSI is found, or error if not present. The way to implement this subfunction can vary, e.g. by keeping in-memory a reverse map of $\mathsf{prmslices}$, or by doing a linear search every time.
Algorithm 6 $\mathtt{ReclaimSlice}(V_{i},\sigma)$
1:$\psi\leftarrow\mathsf{SliceMap}_{i}[\sigma]$
2:$\mathsf{ofld}[\psi]:=\mathsf{free}$
3:$\mathsf{SliceMap}_{i}[\sigma]:=\bot$
4:$\ell\leftarrow\mathtt{ReverseShuffle}(\psi)$
5:if $\ell>\mathsf{octr}$ then: return $\triangleright$ No need to reshuffle in this case.
6:$\mathtt{swap}(\mathsf{prmslices}[\ell],\mathsf{prmslices}[\mathsf{octr}])$
7:$j\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\set{\mathsf{octr},\ldots,\mathsf{mxslc}-1}$triangleright $\mathsf{mxslc}=\lceil\frac{S_{\mathsf{D}}-S_{\mathsf{H}}}{S_{\mathsf{P}}}\rceil$ max number of slices.
8:$\mathtt{swap}(\mathsf{prmslices}[j],\mathsf{prmslices}[\mathsf{octr}])$
9:if $\mathsf{ofld}[\mathsf{prmslices}[\mathsf{octr}]]=\mathsf{free}$ then: $\mathsf{octr}:=\mathsf{octr}-1$
@@ 6.7 Unbounded Number of Volumes
Shufflecake assumes a number $N_{\mathsf{MAX}}$ of possible volumes that can be provided by any device. Even if this limit can be chosen freely by implementations, it would be desirable to have a way for creating unlimited numbers of volumes (subject to space availability) per device. This would not only remove an artificial limitation on the scheme, but would also strengthen its operational security by making any kind of safeword-like technique (as discussed in Section 4.2.5) impossible.
For this to be achievable, volume headers should probably not be adjacent and packed at the beginning of the disk. One idea for further investigation would be to embed every header (except the first, ‘less secret’ one, which is still going to be at the beginning) at random positions within the device, and having them linked by the previous volume header through an ad-hoc pointer field which is always present, and indistinguishable from random without the correct password. Traversing this list of linked headers, however, presents some challenges. In particular, when the user provides one password on volume instantiation, how do we know whether the password is wrong? And, if not, how do we reach the right header unlocked by that password? It might be possible to devise some complex linking scheme for an arbitrary number of “bogus” headers on the device, but in any case the following limitations would apply.
@@
    During $\mathtt{init}$, in case there is a collision during a nxtptr generation over the location of another pre-existing header, the currently generated nxtptr value is discarded and sampled again, until a suitable one is found by brute-forcing (since the total header size is supposed to be negligible in comparison to the device size, this should be very efficient).
@@
The above idea might very well work, but it remains to specify an efficient way to embed this way also the slice maps, which might be larger than one slice. Many options are open to evaluation here, from linked lists starting at the header, to multiple branching pointers.
@@ 6.8 Hidden Shufflecake OS
More concretely, eventually Shufflecake could become itself a full PD-focused Linux distribution, where during installation the user is guided in the process of creating volumes and installing other distributions therein in a guided way. For operational efficiency and security, every OS at layer $j$ should be aware of the filesystem and OS in the volume at layers $i<j$ (which is made possible by the hierarchy among Shufflecake volumes). This would also allow a butler daemon to run from the currently running OS and operate in the background on lower-hierarchy OSes, e.g. by performing system updates, downloading emails, etc., so that all these decoy systems are kept up-to-date even if the user neglects to use them regularly. This would in turn allow to ease the suspicion of an adversary when surrendering a decoy password.
@@
Appendix 0.A Changelog
2023-12-07
First revision of this document. Release reference software Shufflecake v0.4.3.
  • Add this Changelog in appendix.
  • Add Key and Glossary in appendix.
  • Change and restyle symbols and variable names in order to be closer to the reference code implementation.
  • Switch from math-style array indices beginning from $1$ to coding-style array indices beginning from 00.
  • Add paragraph “Handling slice conflicts” in Section 4.1.
  • Add paragraph “Device instantiation” and full algorithm for device instantiation in Section 4.1.
  • Expand Section 6.4 with discussion on block-level corruption recovery (and related Algorithm 5) and on corruption mitigation API.
2023-10-06
First version of this document. Release reference software Shufflecake v0.4.1.
Appendix 0.B Key and Glossary
$B$ Block of data (typically a 4096 bytes string.)
$S_{\mathsf{B}}$ Blocksize, in bits (typically 4096 bytes = 32768 bits).
$\mathsf{device}$ Block device (physical partition, USB drive, file-backed loop device, SD card, etc).
$S_{\mathsf{D}}$ Size (in blocks) of a given device.
$V_{i}$ Logical Shufflecake volume of index $i$ (the higher, the more secret).
$N$ Number of volumes created (from $V_{0}$ to $V_{N-1}$).
$N_{\mathsf{MAX}}$ Maximum number of volumes supported per device (15 in our implementation).
$\mathsf{password}$ User-provided password.
$\mathsf{KEK}$ Key Encryption Key, derived from a password and used to encrypt a DMB cell.
$\mathsf{DMB}$ Device Master Block. Contains $N_{\mathsf{MAX}}$ DMB cells.
$\mathsf{cell}$ DMB cell. Contains (encrypted with a KEK) the VMK for the associated volume.
$\mathsf{VMK}$ Volume Master Key, used to encrypt a VMB.
$\mathsf{VMB}$ Volume Master Block. Contains information on the associated volume.
$S_{\mathsf{H}}$ Size (in blocks) of a device header (DMB + VMBs).
$\mathsf{metadata}$ metadata associated to a volume. Contained in the volume’s VMB.
$\mathsf{VEK}$ Volume Encryption Key, used to encrypt position map and data slices of a volume.
$S_{\mathsf{L}}$ Size (in blocks) of a logical slice.
$S_{\mathsf{P}}$ Size (in blocks) of a physical slice.
$\Delta_{\mathsf{S}}$ Difference (in blocks) between $S_{\mathsf{P}}$ and $S_{\mathsf{L}}$, used to store IVs for CTR mode.
$\psi$ Physical slice index (PSI), i.e., position of a physical slice within a device.
$\sigma$ Logical slice index (LSI), i.e., position of a logical slice within a volume.
$\mathsf{SliceMap}_{i}$ Slice map of volume $V_{i}$.
$\mathsf{numslices}$ Number of logical slices available across all volumes.
$\mathsf{mxslc}$ Maximum supported number of slices for a given device.
$\beta_{\mathsf{L}}$ Logical block address, i.e., index of a block within a volume.
$\beta_{\mathsf{P}}$ Physical block address, i.e., index of a block within a device.
$\mathsf{prmslices}$ Array of permuted PSIs of a device (reshuffled at every instantiation).
$\mathsf{ofld}$ Array of occupation flags, with PSIs as indices and $\mathsf{free}$/$\mathsf{occupied}$ values.
$\mathsf{octr}$ Occupation counter for $\mathsf{prmslices}$: the first $\mathsf{octr}$ elements of $\mathsf{prmslices}$ are $\mathsf{occupied}$.